# Gaussian-LIC: Real-Time Photo-Realistic SLAM with Gaussian Splatting and LiDAR-Inertial-Camera Fusion

### v2

@@ IV-A1 Implementation Details
Our evaluations are run on a desktop PC with an NVIDIA RTX 3090 24 GB GPU, a 3.2GHz Intel Core i7-8700 CPU, and 32 GB RAM. For mapping data preparation, we set $N_{l}$ to 10, $N_{k}$ to 11, and $N_{t}$ to 9. For map initialization, we set $o_{a}$ to 0.1, $o_{b}$ to 0.7, $N_{s}$ to 100,000, and $R$ to 10,000. For map expansion, we set $\tau$ to 0.99. As for map optimization, we set the loss weighting $\lambda$ to 0.2 and the number $K$ of selected keyframes to 100. All the learning rates for Gaussian attributes are kept the same with the vanilla 3DGS [kerbl20233d] but not decayed with schedulers. And we share the same learning rate of the affine matrix with [kerbl2024hierarchical]. Experiments in all tested scenarios share the same hyperparameters for fair comparison.
@@ IV-A2 Datasets
We conduct extensive experiments on three public LiDAR-Inertial-Camera datasets, including the FAST-LIVO dataset [zheng2022fast] and the R3LIVE dataset [lin2022r] with a solid-state LiDAR, and the MCD dataset [mcdviral2024] with a mechanical spinning LiDAR. We test 9 sequences among the three datasets, including (1) hku2 (f0), LiDAR_Degenerate (f1), and Visual_Challenge (f2) from the FAST-LIVO dataset, (2) hku_campus_seq_00 (r0), degenerate_seq_00 (r1), and degenerate_seq_01 (r2) from the R3LIVE dataset, and (3) segments of tuhh_day_02 (m0), tuhh_day_03 (m1), and tuhh_day_04 (m2) sequences in the large-scale MCD dataset [mcdviral2024]. Note that for all tested datasets, we only adopt the left images if the stereo images are provided. The FAST-LIVO dataset and the R3LIVE dataset provide $640\times 512$ images, while the MCD dataset has $640\times 480$ images.
@@ IV-A3 Baselines
Note that when operating in unbounded outdoor environments, depth data from RGB-D cameras with a limited sensing range is often unreliable. Therefore, we first compare our method with the SOTA RGB-only systems, including the NeRF-based approach NeRF-SLAM [rosinol2023nerf] and the 3DGS-based method MonoGS [matsuki2024gaussian]. Additionally, since no open-sourced radiance-field-based LiDAR-Camera SLAM system exists, we adapt the state-of-the-art RGB-D method SplaTAM [keetha2024splatam] to work with LiDAR and monocular images for comparison. Specifically, we merge point clouds from multiple LiDAR scans and obtain dense depth maps through projection, creating pseudo RGB-D images that can be used as input for SplaTAM. For a fair comparison, we adjust its opacity threshold $\tau$ to match ours, allowing more Gaussians in unbounded scenes, which proves to be effective. To evaluate real-time performance, post-processing is disabled for all compared methods.
@@
NeRF-SLAM [rosinol2023nerf] MonoGS [matsuki2024gaussian] SplaTAM [keetha2024splatam] Gaussian-LIC GT
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
@@ IV-B1 Tracking
Based on single-modality sensor input or simple re-rendering loss for tracking, all the compared methods get large drift or even fail in pose estimation across all the tested sequences. On the contrary, our system consistently achieves superior tracking robustness and accuracy by tightly coupling LiDAR-Inertial-Camera data in a continuous-time factor graph [lv2023continuous].
@@ IV-B2 Rendering
We analyze the performance of photo-realistic mapping by checking the quality of RGB rendering results. Metrics including Peak Signal to Noise Ratio (PSNR), Structural Similarity (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) are adopted for analysis [kerbl20233d]. Now that all compared methods badly estimate poses in complex scenarios, we run them with ground truth poses for purely examining the mapping performance. In the FAST-LIVO and R3LIVE datasets where the ground-truth poses are not provided, we run them with our estimated poses instead. Table IV exhibits the rendering results on both the training views and novel views. Qualitative results are also shown in Fig. IV-A3. Lack of precise priors for Gaussian initialization, MonoGS shows limited performance outdoors. With extra depth supervision estimated from DroidSLAM [teed2021droid], NeRF-SLAM displays slightly better results but renders full-resolution images from neural implicit maps. For improved running speed, SplaTAM represents the scene using isotropic Gaussians without considering view-dependent effects, sacrificing visual quality and tending to degenerate in complicated unbounded scenes. With superior geometric priors from LiDAR points and visual SFM points, and the appropriate modeling of sky and camera exposure, our system consistently realizes the best rendering quality across all sequences, no matter on train views or novel views (see Tab IV and Fig. IV-A3)
TABLE II: Runtime (seconds) of different methods on the sequence f0 (span: 105 seconds) with their estimated poses.
@@
NeRF-SLAM [rosinol2023nerf] 105 288 288 25.23
MonoGS [matsuki2024gaussian] 181 387 387 21.42
SplaTAM [keetha2024splatam] 954 1620 2574 17.37
COLMAP [schoenberger2016sfm] + 3DGS [kerbl20233d] - - 6764 32.56
Gaussian-LIC 105 105 105 29.89
Gaussian-LIC w/o acc 105 198 198 29.89
@@ IV-B3 Runtime
We evaluate the real-time performance of all the compared methods, where real-time is defined as completing the processing of data within the period of receiving sensor data, and without extensive post-processing. All approaches rely on their estimated poses rather than privileged ground-truth poses. We record the total runtime of the tracking and mapping modules, including the time spent waiting for and receiving sensory data. For the sequential system SplaTAM, the total runtime is the sum of the time taken by both modules, whereas for the parallel systems, the runtime is determined by the module with the longer time consumption. Tab. II presents the results on sequence f0, which spans 105 seconds. Powered by real-time LIC odometry and carefully designed strategies to accelerate Gaussian map optimization, our method is the only real-time-capable approach while also achieving the highest accuracy. It completes both tracking and mapping immediately after receiving all sensory data. We highlight the importance of these acceleration strategies in Tab. II with the entry Gaussian-LIC w/o acc. Additionally, we show the results for the vanilla 3DGS [kerbl20233d] initialized by COLMAP [schoenberger2016sfm], where the total runtime includes the offline SFM and Gaussian training. Although the offline, batch-based approach with globally consistent poses and extended optimization outperforms online incremental methods, it is unsuitable for real-time robotic applications.
@@ IV-C Ablation Study
Table III presents the results of our ablation study on our proposed method. The removal of any individual component—such as modeling camera exposures, modeling the sky, or incorporating visual SFM points—leads to a decline in rendering quality. As illustrated in Fig. 3, the inclusion of sky Gaussians prevents surface Gaussians from incorrectly modeling the sky, thereby enhancing rendering fidelity. Additionally, integrating visual SFM points helps initialize 3D Gaussians in regions where LiDAR data is unavailable, such as the tops of trees, as shown in Fig. 3.
@@
Figure 3: The significance of sky modeling and incorporating SFM points. (a) and (b) are rendered images on the sequence r0. Without handling the sky, blurring occurs around the buildings because the surface Gaussians are trying to incorrectly fit the sky. (c) LiDAR points (colored red) cannot cover the entire scene in the camera FoV. We alleviate this problem by additionally incorporating visual SFM points (colored green) obtained by online triangulation.
@@ V CONCLUSIONS
In this study, we introduce a real-time, photo-realistic SLAM system that integrates Gaussian Splatting and LiDAR-Inertial-Camera fusion. Unlike previous radiance-field-based SLAM mainly designed for indoor environments with RGB-D or RGB sensors, which struggle with unbounded outdoor scenes and challenging conditions like rapid motions and fluctuating light, our method is built for general use. It enhances pose estimation and mapping quality by additionally fusing LiDAR and IMU data. Our SLAM system accommodates both solid-state and mechanical spinning LiDAR and utilizes visual SFM points to initialize Gaussians in areas not covered by the LiDAR. We’ve also incorporated sky modeling and camera exposure modeling for enhanced rendering. Our system also benefits from optimized strategies for fast Gaussian map optimization. Experiments show its superior performance. In the future, we will try to enhance the accuracy of the odometry by the reconstructed Gaussian map and improve the geometrical reconstruction quality.

### v3

@@ IV-A1 Implementation Details
Our evaluations are conducted on a desktop PC with an NVIDIA RTX 3090 GPU (24 GB VRAM), a 3.2GHz Intel Core i7-8700 CPU, and 32 GB of RAM. For mapping data preparation, we set $N_{l}$ to 10, $N_{k}$ to 11, and $N_{t}$ to 9. For map initialization, we set $o_{a}$ to 0.1, $o_{b}$ to 0.7, $N_{s}$ to $1e5$, and $R$ to $1e4$. For map expansion, we set $\tau$ to 0.99. As for map optimization, we set the loss weighting $\lambda$ to 0.2 and the number $K$ of selected keyframes to 100. All the learning rates for Gaussian attributes are kept the same with the vanilla [2] but not decayed with schedulers. We share the same learning rate of the affine matrix with [54]. Experiments across all tested scenarios use the same hyperparameters to ensure a fair and comprehensive evaluation.
@@ IV-A2 Datasets
We conduct extensive experiments on three public LiDAR-Inertial-Camera datasets, including the FAST-LIVO dataset [38] and the R3LIVE dataset [37] with a solid-state LiDAR, and the MCD dataset [57] with a mechanical spinning LiDAR. We test 9 sequences among the three datasets, including (1) hku2 (f0), LiDAR_Degenerate (f1), and Visual_Challenge (f2) from the FAST-LIVO dataset, (2) hku_campus_seq_00 (r0), degenerate_seq_00 (r1), and degenerate_seq_01 (r2) from the R3LIVE dataset, and (3) segments of tuhh_day_02 (m0), tuhh_day_03 (m1), and tuhh_day_04 (m2) sequences in the large-scale MCD dataset [57]. Note that for all tested datasets, we only adopt the left images if the stereo images are provided. The FAST-LIVO dataset and the R3LIVE dataset provide $640\times 512$ images, while the MCD dataset has $640\times 480$ images.
@@ IV-A3 Baselines
Note that when operating in unbounded outdoor environments, depth data from RGB-D cameras with a limited sensing range is often unreliable. Therefore, we first compare our method with the SOTA RGB-only systems, including the NeRF-based approach NeRF-SLAM [15] and the 3DGS-based method MonoGS [26]. Additionally, since no open-source radiance-field-based LiDAR-Camera SLAM system is currently available, we adapt the state-of-the-art RGB-D method SplaTAM [18] to operate with LiDAR and monocular images for comparison. Specifically, we merge point clouds from multiple LiDAR scans and obtain dense depth maps through projection, creating pseudo RGB-D images that can be used as input for SplaTAM. For a fair comparison, we adjust its opacity threshold $\tau$ to match ours, allowing more Gaussians in unbounded scenes, which proves to be effective. To evaluate real-time performance, post-processing is disabled for all compared methods.
@@
NeRF-SLAM [15] MonoGS [26] SplaTAM [18] Gaussian-LIC GT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ IV-B1 Tracking
Based on single-modality sensor input or simple re-rendering loss for tracking, all the compared methods get significant drift or even fail in pose estimation across all the tested sequences. On the contrary, our system consistently achieves superior tracking robustness and accuracy by tightly coupling LiDAR-Inertial-Camera data in a continuous-time factor graph [39].
@@ IV-B2 Rendering
We analyze the performance of photo-realistic mapping by checking the quality of RGB rendering results. Metrics including Peak Signal to Noise Ratio (PSNR), Structural Similarity (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) are adopted for analysis [2]. Note that all compared methods struggle with pose estimation in complex scenarios. Therefore, we run them with ground-truth poses to purely evaluate their mapping performance. In the FAST-LIVO and R3LIVE datasets, where ground-truth poses are unavailable, we run the compared methods using the more accurate estimated poses from our approach instead. Table I shows the evaluation of rendering results on both the training views and novel views. Qualitative results are also shown in Fig. 2. Lack of precise priors for Gaussian initialization, MonoGS shows degraded performance outdoors. With additional depth supervision estimated from DroidSLAM [30], NeRF-SLAM achieves slightly improved results but at the cost of significantly increased rendering time. For improved running speed, SplaTAM represents the scene using isotropic Gaussians without considering view-dependent effects, sacrificing visual quality and tending to degenerate in complicated unbounded scenes. With superior geometric priors from a combination of LiDAR points and visual SFM points, and the appropriate modeling of sky and camera exposure, our system consistently achieves the best rendering quality across all sequences on the training views. Even more promisingly, the quality of our renderings at novel views surpasses that of other methods, even on their own training views.
TABLE II: Runtime (seconds) of different methods on the sequence f0 (duration: 105 seconds) with their estimated poses.
@@
NeRF-SLAM [15]
105 288 288 25.23
MonoGS [26]
181 387 387 21.42
SplaTAM [18]
954 1620 2574 17.37
COLMAP [58] + 3DGS [2]
- - 6764 32.56
Gaussian-LIC
105 105 105 29.89
Gaussian-LIC w/o acc
105 198 198 29.89
@@ IV-B3 Runtime
We evaluate the real-time performance of all compared methods, defining real-time as completing data processing within the duration of sensor data acquisition, without extensive post-processing. All approaches rely on their estimated poses rather than privileged ground-truth poses. For the sequential system SplaTAM, the total runtime is the sum of the time taken by both tracking and mapping modules, whereas for the parallel systems (NeRF-SLAM, MonoGS, Gaussian-LIC), the runtime is determined by the module with the longer time consumption. Tab. II presents the results on sequence f0 (105s duration). Powered by real-time LIC odometry and carefully designed strategies to accelerate Gaussian map optimization, our method is the only real-time-capable approach while also achieving the highest accuracy. It completes both tracking and mapping immediately after receiving all sensory data. We highlight the importance of the introduced acceleration strategies (see Sec. III-D) in Tab. II with the entry Gaussian-LIC w/o acc. Additionally, we show the results for the vanilla 3DGS [2] initialized by COLMAP [58], where the total runtime includes the offline SFM and Gaussian training. Although the offline, batch-based approach with globally consistent poses and extended optimization outperforms online incremental methods, it is unsuitable for real-time robotic applications.
@@ IV-C Ablation Study
Table III presents the results of the ablation study on our proposed method. The removal of any individual component—such as modeling camera exposures, modeling the sky, or incorporating visual SFM points—leads to a decline in rendering quality. As illustrated in Fig. 3, the inclusion of sky Gaussians prevents surface Gaussians from incorrectly modeling the sky, thereby enhancing rendering fidelity. Additionally, integrating visual SFM points helps initialize 3D Gaussians in regions where LiDAR data is unavailable.
@@
Figure 3: The significance of sky modeling and incorporating SFM points. (a) and (b) are rendered images on the sequence r0. Without handling the sky, blurring occurs around the buildings (top left) because the surface Gaussians are trying to incorrectly fit the sky. (c) LiDAR points (colored red) cannot cover the entire scene in the camera FoV. We alleviate this problem by additionally incorporating visual SFM points (colored green) obtained by online triangulation.
@@ V CONCLUSIONS
In this study, we introduce a real-time, photo-realistic SLAM system that integrates LiDAR-Inertial-Camera odometry and Gaussian Splatting. Unlike previous radiance-field-based SLAM systems, which are primarily designed for indoor scenes with RGB-D or RGB sensors and struggle in unbounded outdoor scenes or under challenging conditions such as rapid motion and fluctuating lighting, our method is specifically developed to overcome these challenges outdoor. It enhances pose estimation and mapping quality by additionally fusing LiDAR and IMU data. Our SLAM system accommodates both solid-state and mechanical spinning LiDAR and utilizes visual SFM points to initialize Gaussians in areas not covered by the LiDAR. We’ve also incorporated sky modeling and camera exposure modeling for enhanced rendering. Our system also benefits from optimized strategies for fast Gaussian map optimization. Experiments show its superior performance. In the future, we will try to enhance the accuracy of the odometry by the reconstructed Gaussian map and improve the geometrical reconstruction quality.
VI ACKNOWLEDGEMENT
This work is supported by the ”Leading Goose” Key R&D Program of Zhejiang Province in China (2025C01069).